OpenStudy (anonymous):

DO NOT READ UNLESS YOU ARE @hartnn Solve the following linear equation for x. 2x + 7 = 2(6 + x) – 4 A. x = -2 B. x = 5/2 @hartnn Can you explain this without giving me the answer? :p

hartnn (hartnn):

elimination method again ?

hartnn (hartnn):

or would you prefer solving it ?

OpenStudy (beccaboo333):

I think others can help you rather than just hartnn.. ._.

OpenStudy (anonymous):

I know @beccaboo333 I just don't want the question to get a ton of answer, cause then I'll get confused. :) @hartnn Which ever way is best. :p I just don't want an answer. I wanna know how to do it.

hartnn (hartnn):

lets try solving it! because you won't always get choices...

hartnn (hartnn):

2(6 + x) = ... know distributive property ?

OpenStudy (anonymous):

LOL! u should change the title or else everyone would be interested to read

hartnn (hartnn):

\(\Large ☺(☻+♥)=☺☻+☺♥\) \(\Large 2(6+x) = ...\)

OpenStudy (anonymous):

26 + 2x?

hartnn (hartnn):

its not just putting numbers together, its multiplication! :) \(2 (6+x) = 2\times 6 + 2 \times x = 12+2x\) got this ?

OpenStudy (anonymous):

yes. :)

hartnn (hartnn):

good! so now we have 2x+7 = 12+2x -4 can you try to simplify this ?? (combine like terms)

OpenStudy (anonymous):

What does X equal?

hartnn (hartnn):

let 'x' be 'x' only its a variable. combine 12 and -4 12-4 = ..?

OpenStudy (anonymous):

-48?

hartnn (hartnn):

12-4 is 8 2x+7 = 8+2x

OpenStudy (anonymous):

Oh! I thought you meant 12x-4! :)

hartnn (hartnn):

now subtract 2x from both sides! \(2x-2x+7 = 8+2x-2x\)

hartnn (hartnn):

and 2x-2x = (2-2)x = 0

hartnn (hartnn):

so what are you left with ?

OpenStudy (anonymous):

2?

hartnn (hartnn):

\(2x-2x+7 = 8+2x-2x \\ 0+7 = 8+0 \\ 7=8 \) is this even possible ?

OpenStudy (anonymous):

Oh! Typo sorry. :) My key board gets jammed. :) 5/2

hartnn (hartnn):

not 5/2 is "no solution" one of the choices ?

OpenStudy (anonymous):

Oh, yeah I didn't even see that. :)

hartnn (hartnn):

thats because we got 7=8 which is not possible for any value of x so no solutions

OpenStudy (anonymous):

I see. :)